OpenStudy (anonymous):

The table below represents the displacement of a duck from its nest as a function of time. time(x) displacement(y) 0 7 1 29 2 51 3 73 4 95 Part A: What is the y-intercept of the function and what does this tell you about the duck? (4 points) Part B: Calculate the average rate of change of the function represented by the table between x=2 to x=4 hours? (4 points) Part C: What would be the domain of the function if the duck continued to walk at this rate until it traveled 227 miles from the

OpenStudy (anonymous):

Part A: What is the y-intercept? What does this tell you about the duck? Part B: What is the average rate of change from x = 2 to x = 4? Show your work. Part C: Write a compound inequality of the domain if he traveled 227 miles from the nest?

OpenStudy (anonymous):

@aum @perl

OpenStudy (aum):

y-intercept is the y-value when x = 0. From the table, what is the y-value when x = 0?

OpenStudy (anonymous):

7

OpenStudy (aum):

Correct. y-intercept = 7 What does this tell you about the duck? At time x = 0, the duck is 7 units away from the nest. That means the duck was NOT in its nest at time x = 0.

OpenStudy (aum):

Part B: What is the average rate of change from x = 2 to x = 4? Show your work. Average rate of change = \(\Large \frac{y_2-y_1}{x_2-x_1}\) Read the y-values from the table when x = 2 and when x = 4 Average rate of change = \( \Large \frac{95-51}{4-2} = ?\)

OpenStudy (anonymous):

44/2

OpenStudy (anonymous):

@aum

OpenStudy (aum):

44 / 2 = 22. Part B: What is the average rate of change from x = 2 to x = 4? 22 Show your work: See previous reply

OpenStudy (aum):

C) Slope = 22 (same as average rate of change found in part B.) y = mx + b y = 22x + b when x = 0, y = 0 + b y = b From the table, when x = 0, y = 7. Therefore, b = 7 y = 22x + 7 Part C: Write a compound inequality of the domain if he traveled 227 miles from the nest? Set y = 227 in y = 22x + 7 227 = 22x + 7 227 - 7 = 22x 22x = 220 x = 10 Domain is x = 0 to x = 10. Or, domain is: [0, 10].

OpenStudy (aum):

Compound inequality for the domain: \(\large 0 \le x \le 10\)